OpenStudy (lisa2345):

Need help on test for Elimination Method questions below just give answers in order need them asap limited time!! Reward will be given. :) 1. 3x-y=14 -2x-y=-1 2. 2x-3y=5 x+2y=-1 3. -3x+4y=29 3x+2y=-17 4. y=7x+4 x+y=4 5. y=4x+2 y=6x-10

OpenStudy (lisa2345):

@Compassionate @Directrix @dumbcow @triciaal

OpenStudy (triciaal):

1)subract eqn 2 from 1 solve for x 2) multiply eqn 2 by 2 subtract solve for y

OpenStudy (triciaal):

3)just add 6y = 22 y = then find x

OpenStudy (triciaal):

4.)eqn 2 y = 4-x eqn subtract eqns solve for x

OpenStudy (triciaal):

5 set = 12 = 2x x = 6 y = 6*6 - 10 = 26

OpenStudy (lisa2345):

This is still hard to understand

OpenStudy (triciaal):

did you try what I suggested?